Four people are pushing a body with forces:

310N in the direction 045

280N in the direction 135

400N in the direction 80. Calculate sum of vertical and sum of horizontal component respectively.?


1
Expert's answer
2022-02-19T04:56:13-0500

"F_1" = 310 N,

"F_2" = 280 N,

"F_3" = 400 N.

Sum of vertical component:

"F_y=F_{1y}+F_{2y}+F_{3y}=F_1sin45\degree+F_2sin45\degree+F_3sin80\degree=\newline 
310\cdot 0.7+280\cdot 0.7+400\cdot 0.98=811.1\space N"

Sum of horizontal component:

"F_x=F_{1x}-F_{2x}+F_{3x}=F_1cos45\degree-F_2cos45\degree+F_3cos80\degree=\newline 
310\cdot 0.7-280\cdot 0.7+400\cdot 0.17=90.7\space N"